# Mildly Accurate Computationally Differentially Private Inner Product Protocols Imply Oblivious Transfer

### v1

@@ C.0.0.3 $\mathsf{A}$’s privacy:
in contradiction to Equation 31. Next, by Appendix C there exists an algorithm $\mathsf{D}^{\prime}$ such that
@@
Appendix D Missing Proofs
[…]
D.1 Proving Section D.3
We make use of the following fact.
[…]
We now prove Section D.3, restated below.
[…]
Proof.
(denote this set by ${\cal I}$). Thus, we deduce for $I\leftarrow{\cal I}$ that
$\geq 1/\sqrt{m}.$
On the other hand, note that
where the second inequality holds by Section D.3. Thus, we conclude that
$<1/\sqrt{m},$
D.2 Proving Sections 3.3 and 3.3
@@ Appendix D Missing Proofs
Fix $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ and $i\in[n]$. By Section D.2, it is sufficient to prove that
@@
Since the above holds for any $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$, we conclude by Section D.2 that
@@
The proof of Section D.3 uses the following fact.
@@
We first prove Section D.3 using Section D.3.
Proof of Section D.3.
@@
where the second inequality holds by Section D.3. Thus, we conclude that
@@
Using Section D.3, we now prove Section 5.1, restated below.
@@
On the other hand, for every fixing of $(z,w)\in\operatorname{Supp}(Z,W)$, we can apply Section D.3 with the function $F_{z,w}(r)=F(r,z_{r},w)$ and with $\alpha=\frac{\gamma}{16}$ to obtain that
@@
We next prove Equations 36 and 37 using Equations 39 and 40.
@@
We next prove the lemma using Equations 41 and 42.
@@
It is left to prove Equation 37. Observe that Equations 41 and 42 imply with probability at least $\gamma/2-\frac{512}{n\gamma^{2}}$ over $t=(i,z,w)\in(I,Z,W)$, we have $\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|\geq\frac{\gamma}{2}$. Therefore, conditioned on event $E_{t}$ (i.e., $\forall x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}:\>\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\leq\frac{\gamma}{16}$), we have for such triplets $t=(i,z,w)$ that

### v2

@@ B.0.0.3 $\mathsf{A}$’s privacy:
in contradiction to Equation 30. Next, by Appendix B there exists an algorithm $\mathsf{D}^{\prime}$ such that
@@
Appendix C Missing Proofs
C.1 Proving Sections 4.3 and 4.3
@@ Appendix C Missing Proofs
Fix $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ and $i\in[n]$. By Section C.1, it is sufficient to prove that
@@
Since the above holds for any $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$, we conclude by Section C.1 that
@@
The proof of Section C.2 uses the following fact.
@@
We first prove Section C.2 using Section C.2.
Proof of Section C.2.
@@
where the second inequality holds by Section C.2. Thus, we conclude that
@@
Using Section C.2, we now prove Section 6.1, restated below.
@@
On the other hand, for every fixing of $(z,w)\in\operatorname{Supp}(Z,W)$, we can apply Section C.2 with the function $F_{z,w}(r)=F(r,z_{r},w)$ and with $\alpha=\frac{\gamma}{16}$ to obtain that
@@
We next prove Equations 34 and 35 using Equations 37 and 38.
@@
We next prove the lemma using Equations 39 and 40.
@@
It is left to prove Equation 35. Observe that Equations 39 and 40 imply with probability at least $\gamma/2-\frac{512}{n\gamma^{2}}$ over $t=(i,z,w)\in(I,Z,W)$, we have $\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|\geq\frac{\gamma}{2}$. Therefore, conditioned on event $E_{t}$ (i.e., $\forall x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}:\>\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\leq\frac{\gamma}{16}$), we have for such triplets $t=(i,z,w)$ that